# Unveiling User Preferences: A Knowledge Graph and LLM-Driven Approach for Conversational Recommendation

### v1

@@ Abstract.
Conversational Recommender Systems (CRSs) aim to provide personalized recommendations through dynamically capturing user preferences in interactive conversations. Conventional CRSs often extract user preferences as hidden representations, which are criticized for their lack of interpretability. This diminishes the transparency and trustworthiness of the recommendation process. Recent works have explored combining the impressive capabilities of Large Language Models (LLMs) with the domain-specific knowledge of Knowledge Graphs (KGs) to generate human-understandable recommendation explanations. Despite these efforts, the integration of LLMs and KGs for CRSs remains challenging due to the modality gap between unstructured dialogues and structured KGs. Moreover, LLMs pre-trained on large-scale corpora may not be well-suited for analyzing user preferences, which require domain-specific knowledge. In this paper, we propose COMPASS (Compact Preference Analyzer and Summarization System), a plug-and-play framework that synergizes LLMs and KGs to unveil user preferences, enhancing the performance and explainability of existing CRSs. To address integration challenges, COMPASS employs a two-stage training approach: first, it bridges the gap between the structured KG and natural language through an innovative graph entity captioning pre-training mechanism. This enables the LLM to transform KG entities into concise natural language descriptions, allowing them to comprehend domain-specific knowledge. Following, COMPASS optimizes user preference modeling via knowledge-aware instruction fine-tuning, where the LLM learns to reason and summarize user preferences from both dialogue histories and KG-augmented context. This enables COMPASS to perform knowledge-aware reasoning and generate comprehensive and interpretable user preferences that can seamlessly integrate with existing CRS models for improving recommendation performance and explainability. Our experiments on benchmark datasets demonstrate the effectiveness of COMPASS in improving various CRS models.
Conversational Recommender System; Large Language Model; Knowledge Graph
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
1. Introduction
Figure 1. Conventional CRSs lack explainability in user preference modeling by extracting hidden representation, while our approach enhances transparency by generating interpretable user preference in text.
@@ 1. Introduction
Providing personalized, context-aware recommendations that align with users’ changing preferences and situational needs is a critical challenge across various domains such as e-commerce, streaming platforms, and other online services. In recent years, conversational recommender systems (CRSs) have emerged as a promising approach, harnessing the power of natural language interactions to unravel user preferences (Jannach et al., 2021; Gao et al., 2021). By engaging the user in interactive conversations, CRSs enable a better understanding of the user’s evolving interests and guide them toward products, services, and information that best meet their immediate requirements (Christakopoulou et al., 2016; Li et al., 2018; Liu et al., 2023b). Despite the success, CRSs face unique challenges in precisely modeling user preferences from the semantically rich and dynamically evolving dialogues, due to the challenge of natural language understanding (Jannach et al., 2021). CRSs must capture user’s intent and interest from their natural language inputs, which are often ambiguous and context-dependent (Li et al., 2018; Zhou et al., 2020). Moreover, as the conversation progresses, CRSs need to continuously update and refine their understanding of user preferences in real-time (Sun and Zhang, 2018; Friedman et al., 2023). As the user reveals more about their preferences, CRSs must dynamically adapt their preference modeling to integrate the user’s immediate interests expressed in recent interactions with their overall preferences developed over time.
@@
Conventional CRSs mainly rely on item-centric approaches (Chen et al., 2019; Zhou et al., 2020), focusing on items mentioned during the conversation to model user preferences. Recent studies have incorporated pre-trained language models (PLMs) to enhance both natural language understanding and user preference modeling (Wang et al., 2022c, a). However, these methods often fail to infer implicit preferences or reason about underlying motivations beyond explicit item mentions, leading to a superficial understanding of user intent. Moreover, they typically represent user preferences as hidden embeddings, leaving it unclear what specific preference is being considered when making a recommendation, as shown at the top of Figure 1. This ambiguity makes it challenging to verify the underlying reasons and results behind the recommendations, hindering the transparency and accountability of the recommender system.
@@
Recent advancements in Large Language Models (LLMs), which significantly scale up PLMs in terms of parameter size and pre-training data, have demonstrated exceptional capabilities in both natural language processing (Zhao et al., 2023; Jiang et al., 2023; Dubey et al., 2024) and reasoning (Wei et al., 2023; Wang et al., 2024). These models have shown promise in generating natural language explanations, particularly in recommendation systems where explainability is critical. Recent works like X-REC (Ma et al., 2024) and SLIM (Wang et al., 2024) have demonstrated the LLM’s ability to reason over user historical behaviors and generate interpretable recommendation explanations, enhancing both explainability and performance. Nevertheless, these methods only focus on the traditional recommendation settings, where user preferences are static and inferred from their historical behaviors. CRSs face unique challenges, necessitating real-time preference modeling and recommendations that adapt to users’ changing preferences during ongoing conversations. Although LLMs have shown promise in CRS tasks such as evaluation (Wang et al., 2023b; Zhu et al., 2024), zero-shot recommendation capabilities (He et al., 2023), and task planning (Feng et al., 2023; Fang et al., 2024; Li et al., 2024), the role of explainable user preferences in LLM-based CRSs still remains unexplored.
While LLMs have shown promise in CRS tasks, they still face limitations in incorporating domain-specific knowledge and keeping up-to-date with item information, which are essential for accurate preference modeling and recommendations. Knowledge Graphs (KGs) have proven effective in addressing these limitations (Chen et al., 2019; Zhou et al., 2020; Wang et al., 2022c), as they provide a rich, structured context of items and their relationships, offering domain-specific insights that enable more accurate and explainable recommendations. Despite their demonstrated potential, integrating the structured knowledge of KGs with the reasoning and language capabilities of LLMs presents a new set of challenges:
@@
  • Modality Gap (Challenge 1): There exists a significant modality gap between KGs and LLMs, which hinders the LLM’s ability to understand and interpret KG information. While LLMs process sequences of tokens that represent natural language, KGs represent information in a structured, graph-based format. This difference makes it difficult for LLMs to directly interpret the entities and relationships encoded in KGs, limiting their ability to leverage domain knowledge for preference modeling.
@@
  • Cross-Modal Reasoning (Challenge 2): Effectively reasoning over both KG information and conversation data to infer user preferences is a complex task. LLMs, despite their strong natural language processing capabilities, are not inherently designed to perform this cross-modal reasoning. Therefore, they face difficulties in analyzing and synthesizing insights from graph-structured knowledge alongside dialogue history. This limitation hinders their ability to identify relevant patterns across both sources and to perform the knowledge-aware reasoning necessary for comprehensive user preference modeling, in which both domain knowledge and user interactions are essential.
@@
To address the above challenges, we propose Compact Preference Analyzer and Summarization System (COMPASS), a novel framework that synergizes LLMs and KGs to unveil user preferences, improving both recommendation performance and the explainability of existing CRS. COMPASS directly tackles the limitations of current CRSs and the challenges of integrating LLMs with KGs through a two-stage process. First, we introduce a graph entity captioning pre-training mechanism that transforms KG structures into natural language descriptions. This allows the LLM to comprehend domain-specific information and bridge the modality gap (Challenge 1). We leverage a Graph Neural Network (GNN) to capture structural information from the KG and represent it as entity embeddings. These embeddings are then input into the LLM to produce textual descriptions of the entities along with relevant details from their neighbors. In this way, we align graph-structured knowledge with natural language, enabling the LLM to better interpret KG information. Building upon this alignment, COMPASS employs knowledge-aware instruction fine-tuning to improve the LLM’s ability to reason about user preferences from dialogue histories and KG-augmented contexts. These KG-augmented contexts consist of relevant entity information and relationships extracted from the KG, providing a rich background for inference beyond the conversation history alone. Through carefully designed instructions, we enhance the LLM’s capability to perform cross-modal reasoning (Challenge 2) by analyzing conversation history and cross-referencing with KG-augmented information. This instruction tuning process enhances the LLM’s ability to extract explicit mentions, infer implicit interests, and reason about preferences in relation to various item attributes. Consequently, as shown in the bottom of Figure 1, COMPASS generates comprehensive and interpretable user preferences in text that capture both overall preferences and current interests. To leverage these insights, we introduce an adaptive gating mechanism that integrates summarized preferences into existing CRS models, boosting recommendation performance and explainability without requiring architectural changes. Our main contributions can be summarized as follows:
@@
Figure 2. The overall framework of our COMPASS. COMPASS consists of three components: the graph encoder, the adapter, and the LLM. The adapter aligns the knowledge graph to the LLM. COMPASS follows a two-stage training paradigm - (a) Graph Entity Captioning and (b) Knowledge-aware Instruction Tuning. Once trained, COMPASS can be easily integrated with existing CRS models for (c) Preference Enhanced Recommendation.
[…]
The overall framework of our COMPASS.
2. Preliminaries
Conversational Recommendation. Let $i$ denote a candidate item from the set of items $\mathcal{I}$, and let $w$ denote a word in the vocabulary $\mathcal{V}$. A dialogue $\mathcal{D}$ between a user and the recommender system consists of a sequence of utterances $\mathcal{D}=[u_{1},\ldots,u_{T}]$, where $u_{t}=[w_{1}^{t},\ldots,w_{m}^{t}]$ is the $t$-th utterance composed of $m$ words, and $T$ is the maximum number of turns in the dialogue. As the conversation progresses, the dialogue history up to the turn $t$ is denoted as $H_{t}=[u_{1}:u_{t}]$, where $[u_{1}:u_{t}]$ signifies the chronological sequence of utterances from the first to the $t$-th turn. The CRS estimates the user’s preferences based on $H_{t}$, them recommend $K$ items from $\mathcal{I}$, which are used to generate the next utterance $u_{t+1}$. Note that $\mathcal{I}_{t}$ can be empty when no recommendation is needed. In such cases, the CRS may raise a clarification question or generate a casual conversation response.
@@ 2. Preliminaries
(1) $\mathcal{P}_{t}=f(I_{p},H_{t},\mathcal{E}_{t}^{m},\mathcal{G}),$
@@
(2) $\mathcal{I}_{t}=f_{\text{crs}}(g(\mathcal{P}_{t}),H_{t},\mathcal{G}),$
@@
3. Methodology
We present an overview of COMPASS, followed by detailed descriptions of each component and the training process.
3.1. Overview
@@
3.2. Model Architecture
3.2.1. Graph Encoder
In COMPASS, the KG is a crucial source of domain-specific information that provides extra context for understanding attributes and relationships of items. To efficiently convert the structured knowledge in the KG into a format understandable by the LLM for preference analysis and summarization, we utilize a Relational Graph Convolutional Network (R-GCN) (Schlichtkrull et al., 2018) to capture the complex graph structure and generate entity embeddings. The R-GCN is well-suited for modeling KGs due to its ability to handle multi-relational data and capture higher-order dependencies between entities.
To initialize the entity embeddings, we leverage the textual descriptions $\mathcal{X}$ associated with the entities in the KG. Specifically, for each entity $e\in\mathcal{E}$, we encode its description $x_{e}$ using a pre-trained language model (PLM). This provides a rich semantic foundation for the graph learning process. The R-GCN then captures both entity-level information and the overall graph structure through iterative message passing, which is particularly important for understanding the relationships between items and attributes. Formally, the representation of an entity $e$ at the $l$-th layer is calculated as follows:
(3) $\mathbf{h}_{e}^{(0)}=\text{PLM}(x_{e}),$
(4) $\mathbf{h}_{e}^{(l+1)}=\sigma\left(\sum_{r\in\mathcal{R}}\sum_{e^{\prime}\in{\mathcal{E}_{e}^{r}}}\frac{1}{Z_{e,r}}\mathbf{W}_{r}^{(l)}\mathbf{h}_{e^{\prime}}^{(l)}+\mathbf{W}_{e}^{(l)}\mathbf{h}_{e}^{(l)}\right),$
@@ 3.2.1. Graph Encoder
3.2.2. Graph-to-Text Adapter
The entity embeddings produced by the graph encoder, while rich in structural information, exist in a different representational space from the LLM’s textual token representations. This makes it challenging for the LLM to effectively reason with the KG-augmented context to generate user preferences. To bridge this semantic gap and enable effective knowledge integration, we introduce an adapter module that creates a mapping between graph-structured entity embeddings and the LLM’s textual domain. Specifically, the adaptation process is defined as:
(5) $\mathbf{h}_{e}^{\tau}=f_{P}(\mathbf{h}_{e}),$

### v2

# Reasoning over User Preferences: Knowledge Graph‑Augmented LLMs for Explainable Conversational Recommendations

@@ Abstract
Conversational Recommender Systems (CRSs) aim to provide personalized recommendations by capturing user preferences through interactive dialogues. Explainability in CRSs is crucial as it enables users to understand the reasoning behind recommendations, increasing system transparency and trustworthiness. However, current CRSs often leverage knowledge graphs (KGs) or language models to extract and represent user preferences as latent vectors, which limits their explainability. Large language models (LLMs) offer powerful reasoning capabilities that can bridge this gap by generating human-understandable preference summaries. However, effectively reasoning over user preferences in CRSs remains challenging as LLMs pre-trained on large-scale corpora may not be well-suited for analyzing user preferences, which requires domain-specific knowledge. While KGs provide rich domain knowledge, integrating them with LLMs encounters a significant modality gap between structured KG information and unstructured conversations. In this paper, we propose COMPASS (Compact Preference Analyzer and Summarization System), a plug-and-play framework that synergizes LLMs and KGs to reason over user preferences, enhancing the performance and explainability of existing CRSs. COMPASS employs a two-stage training approach: first, it bridges the gap between the structured KG and natural language through novel graph entity captioning pre-training. This enables the LLM to transform KG entities into natural language descriptions, allowing it to comprehend domain-specific knowledge. Next, COMPASS optimizes user preference reasoning via knowledge-aware instruction fine-tuning, where the LLM learns to reason and summarize user preferences from dialogue histories and KG-augmented context. This enables COMPASS to perform knowledge-aware reasoning and generate interpretable user preferences that can seamlessly integrate with existing CRS models for improving recommendation performance and explainability. Our experiments on benchmark datasets demonstrate the effectiveness of COMPASS in improving various CRS models.
11footnotetext: Corresponding author.
I Introduction
@@
Providing personalized, context-aware recommendations that align with users’ changing preferences and situational needs is a critical challenge across various domains such as e-commerce, streaming platforms, and other online services. In recent years, conversational recommender systems (CRSs) have emerged as a promising approach, harnessing the power of natural language interactions to unravel user preferences [1, 2]. By engaging the user in interactive conversations, CRSs enable a better understanding of the user’s evolving interests and guide them toward products, services, and information that best meet their immediate requirements [3, 4].
Despite the success, CRSs face unique challenges in precisely modeling user preferences from semantically rich and dynamically evolving dialogue. They must interpret often ambiguous and context-dependent natural language inputs [4, 5], while continuously updating and refining their understanding of user preferences in real-time [6], integrating immediate interests with overall preferences developed over time.
Figure 1: Conventional CRSs lack explainability in user preference modeling by extracting hidden representation, while our approach enhances transparency by generating interpretable user preference in text.
@@
Conventional CRSs mainly depends on item-centric approaches [7, 5], focusing on explicitly mentioned items in the conversation to model user preference. Recent studies incorporating pre-trained language models (PLMs) to enhance both natural language understanding and user preference modeling [8, 9]. However, these methods often fail to infer implicit preferences or reason about underlying motivations beyond explicit item mentions, leading to a superficial understanding of user intent. Moreover, they typically represent user preferences as hidden embeddings, leaving it unclear what specific preference is being considered when making a recommendation, as shown at the top of Figure 1. This ambiguity makes it challenging to verify the underlying reasons and results behind the recommendations, hindering the transparency and accountability of the recommender system.
@@
Recent advancements in large language models (LLMs), have demonstrated exceptional capabilities in both language understanding [10, 11] and complex reasoning [12]. These capabilities make LLMs promising for explainable recommendations, and recent works [13, 14, 15] have demonstrated the LLM’s ability to reason over user historical behaviors and generate interpretable recommendation explanations. Nevertheless, these methods only focus on the traditional recommendation settings, where user preferences are static. In contrast, CRSs require more sophisticated reasoning to model user preferences that evolve dynamically throughout conversation. While recent work has explored LLMs for various CRS tasks, including evaluation [16, 17], zero-shot recommendation [18], and task planning [19, 20], the role of explainable user preferences in LLM-based CRSs remains unexplored.
Enabling LLMs to reason over user preferences requires more than their general language capabilities. Effective preference reasoning requires domain-specific knowledge about item attributes, relationships, and up-to-date information that LLMs inherently lack. Knowledge graphs (KGs) provide this essential structured knowledge through rich semantic relationship and have proven effective in CRSs [7, 5, 8, 21]. Despite their demonstrated potential, integrating KGs with LLMs for preference reasoning presents fundamental challenges that existing approaches have not addressed:
@@
  • Modality Gap (Challenge 1): There exists a significant modality gap between KGs and LLMs. While LLMs process sequences of tokens representing natural language, KGs represent information in structured, graph-based formats with explicit entity relationships. This representational mismatch limits LLMs’s ability to directly interpret and reason with the domain knowledge encoded in KGs.
@@
  • Cross Modal Reasoning (Challenge 2): Effectively reasoning over both the KG and conversation to infer user preferences is a complex task. LLMs, despite their strong natural language processing capabilities, are not inherently designed to perform this cross-modal reasoning. Therefore, they face difficulties in analyzing and synthesizing insights from graph-structured knowledge alongside dialogue history. This limitation hinders their ability to identify relevant patterns across both sources and to perform the knowledge-aware reasoning necessary for comprehensive user preference modeling.
@@
To address the above challenges, we propose Compact Preference Analyzer and Summarization System (COMPASS), a novel framework that augments the LLM with the KG to enable explainable preference reasoning, improving both recommendation performance and the explainability of existing CRSs. COMPASS introduces a two-stage training paradigm that fundamentally transforms how LLMs process and reason with structured knowledge for preference analysis. First, we introduce graph entity captioning pre-training that transforms KG structures into natural language descriptions. This allows the LLM to comprehend domain-specific information and bridge the modality gap (Challenge 1). We employ a Graph Neural Network (GNN) to encode structural information from the KG as entity embeddings, which are then projected into the LLM’s semantic space via an adapter module. The LLM learns to decode these projected embeddings into textual descriptions of entities and their relationships, effectively translating graph-structured knowledge into a format compatible with its reasoning mechanisms. Building upon this alignment, COMPASS employs knowledge-aware instruction fine-tuning to enhance the LLM’s ability to reason about user preferences from dialogue histories and KG-augmented contexts. These KG-augmented contexts consist of relevant entity information and relationships extracted from the KG, providing a rich background for inference beyond the conversation history alone. Through carefully designed instructions, we enhance the LLM’s capability to perform cross-modal reasoning (Challenge 2) by analyzing conversation history and cross-referencing with KG-augmented information. This instruction tuning process enhances the LLM’s ability to extract explicit mentions, infer implicit interests, and reason about preferences in relation to various item attributes. Consequently, as shown in the bottom of Figure 1, COMPASS generates comprehensive and interpretable user preferences in text that capture both overall preferences and current interests. To leverage these insights, we introduce an adaptive gating mechanism that integrates summarized preferences into existing CRSs, boosting recommendation performance and explainability without requiring architectural changes. Our main contributions can be summarized as follows:
@@
II Related Work
Knowledge graphs enable systems to infer implicit user interests and capture item relationships [7, 5], Conversational Recommender System. Conversational recommender systems (CRSs) understand user preferences through interactive dialogue to provide personalized recommendations [1]. These systems face a unique challenge in modeling user preferences from typically brief and sparse conversational data. To address this, recent works incorporate external knowledge sources into CRSs to enhance preference modeling. Knowledge graphs enable systems to infer implicit user interests and capture item relationships [7, 5],, while user reviews provide additional insights into item attributes and user opinions [22, 23], helping to address the challenge of limited contextual information. Moreover, pre-trained language models like DialoGPT [8, 9] and BART [24, 21] has further improved contextual understanding and language generation capabilities in CRSs.
More recently, the advent of powerful LLMs has introduced new opportunities for CRSs. Studies have focused on improving LLM evaluation through enhanced user simulators [16, 17], exploring zero-shot recommendation capabilities [18], using graph-based retrieval augmentation for in-context learning [25], and utilizing LLMs for task planning [19, 20]. However, one critical aspect that remains underexplored is the role of explainable user preference in CRSs. Although some conceptual designs propose using LLMs to manage user preferences [26], they lack quantitative validation. MemoCRS [27] addresses user preference memories in sequential CRS but does not fully harness LLM’s reasoning abilities for preference generation. Our work introduces the first framework that augments LLMs with KG for reasoning over user preferences in CRSs, enhancing both preference modeling and recommendation explainability. Our work introduces the first framework that augments LLMs with KG for reasoning over user preferences in CRSs, enhancing both preference modeling and recommendation explainability.
Explainable Recommendation. Explainable recommendation aims to enhance user trust and decision-making by providing transparent rationales for recommendations. Early research leveraged user-item interactions and content features to generate simple explanations [28], using techniques like recurrent neural networks [29], attention mechanisms [30], and graph neural networks [31]. Recent developments in LLMs have opened new avenues for explainable recommendation. Recent LLMs have shown potential in generating explanations by analyzing user interactions [13, 14], but primarily operate within static user profiles where preferences are inferred from historical behavior. Our work introduces a novel KG-augmented LLM framework for CRSs, enhancing both preference modeling and recommendation explainability in dynamic interaction contexts.
III Preliminaries
Figure 2: The overall framework of our COMPASS. COMPASS consists of three components: the graph encoder, the adapter, and the LLM. The adapter aligns the knowledge graph to the LLM. COMPASS follows a two-stage training paradigm - (a) Graph Entity Captioning and (b) Knowledge-aware Instruction Tuning. Once trained, COMPASS can be easily integrated with existing CRS models for (c) Preference Enhanced Recommendation.
Conversational Recommendation. Let $i$ denote a candidate item from the set of items $\mathcal{I}$, and let $w$ denote a word in the vocabulary $\mathcal{V}$. A dialogue $\mathcal{D}$ between a user and the recommender system consists of a sequence of utterances $\mathcal{D}=[u_{1},\ldots,u_{T}]$, where $u_{t}=[w_{1}^{t},\ldots,w_{m}^{t}]$ is the $t$-th utterance composed of $m$ words, and $T$ is the maximum number of turns in the dialogue. As the conversation progresses, the dialogue history up to the turn $t$ is denoted as $H_{t}=[u_{1}:u_{t}]$, where $u_{1}:u_{t}]$ signifies the chronological sequence of utterances from the first to the $t$-th turn. The CRS estimates the user’s preferences based on $H_{t}$, then recommend $K$ items from $\mathcal{I}$, which are used to generate the next utterance $u_{t+1}$. Note that $\mathcal{I}_{t}$ can be empty when no recommendation is needed. In such cases, the CRS may raise a clarification question or generate a casual conversation response.
@@ III Preliminaries
$\mathcal{P}_{t}=f(I_{p},H_{t},\mathcal{E}_{t}^{m},\mathcal{G}),$ (1)
@@
$\mathcal{I}_{t}=f_{\text{crs}}(g(\mathcal{P}_{t}),H_{t},\mathcal{G}),$ (2)
@@
IV Methodology
IV-A Overview
@@
IV-B Model Architecture
IV-B1 Graph Encoder
In COMPASS, the KG is a crucial source of domain-specific information that provides extra context for understanding attributes and relationships of items. To efficiently convert the structured knowledge in the KG into a format understandable by the LLM for preference reasoning and summarization, we utilize a Relational Graph Convolutional Network (R-GCN) [32] to capture the complex graph structure and generate entity embeddings. The R-GCN is well-suited for modeling KGs due to its ability to handle multi-relational data and capture higher-order dependencies between entities.
To initialize the entity embeddings, we leverage the textual descriptions $\mathcal{X}$ associated with the entities in the KG. Specifically, for each entity $e\in\mathcal{E}$, we encode its description $x_{e}$ using a PLM. This provides a rich semantic foundation for the graph learning process. The R-GCN then captures both entity-level information and the overall graph structure through iterative message passing, which is particularly important for understanding the relationships between items and attributes. Formally, the representation of an entity $e$ at the $l$-th layer is calculated as follows:
$\mathbf{h}_{e}^{(0)}=\text{PLM}(x_{e}),$ (3)
$\mathbf{h}_{e}^{(l+1)}=\sigma\left(\sum_{r\in\mathcal{R}}\sum_{e^{\prime}\in{\mathcal{E}_{e}^{r}}}\frac{1}{Z_{e,r}}\mathbf{W}_{r}^{(l)}\mathbf{h}_{e^{\prime}}^{(l)}+\mathbf{W}_{e}^{(l)}\mathbf{h}_{e}^{(l)}\right),$ (4)
@@ IV-B1 Graph Encoder
IV-B2 Graph-to-Text Adapter
The entity embeddings from the graph encoder exist in a different representational space from the LLM’s textual token representations. This makes it challenging for the LLM to effectively reason with the KG-augmented context to generate user preferences. To bridge this semantic gap and enable effective knowledge integration, we introduce an adapter module that creates a mapping between graph-structured entity embeddings and the LLM’s textual domain. Specifically, the adaptation process is defined as:
$\mathbf{h}_{e}^{\tau}=f_{P}(\mathbf{h}_{e}),$ (5)